OpenStudy (anonymous):

What are the solutions to the equation lx-3l = 14 ? a. x=-11 or 17 b. x= -11 or x=11 C. x= 17 only d. x= 11 only

OpenStudy (anonymous):

the absolute value of what number is 14?

OpenStudy (nurali):

lx-3l = 14 x-3=14 or x-3=-14 x=14+3 or x=-14+3 x=17 or x=-11

OpenStudy (anonymous):

Nurali... read the code of conduct

OpenStudy (anonymous):

code of conduct?